# Supplemetary Material for the paper

### v1

@@ Appendix C Supporting Lemmas and the Proof of the Main Theorem
Lemma C.2.
Let the domain of interest be $\mathcal{D}:=\{(\mathbf{g}(t),\mathbf{V}(t)):|g_{n}(t)|\leq B_{g},|V_{nl}(t)|\leq B_{V}\;\forall n\in[N],l\in[L]\}$ for some constants $B_{g}>0$ and $B_{V}>0$. For the system evolving under continuous gradient flow, the loss function $\mathcal{L}(\mathbf{g}(t),\mathbf{V}(t))=\|\mathbf{Y}-\mathbf{A}(\mathbf{g}(t)^{\odot 2}\mathbf{1}_{L}\odot\mathbf{V}(t))\|_{F}^{2}$ is $\beta$-smooth over $\mathcal{D}$, where $\mathbf{A}$ is a $\mu$-coherent matrix with $l_{2}$-normalized columns. The Lipschitz smoothness constant $\beta$ is given by:
$\beta=16NL^{3/2}\left((N+1)\mu C^{4}+MB_{Y}C\right),$ (143)
where $C\geq\max\{1,B_{g},2B_{V}\}$ and $B_{Y}=\underset{n\in[N],\ l\in[L]}{\max}\{Y_{nl}\}$.
Proof.
For brevity, we omit the explicit time dependency $(t)$ for variables $\mathbf{g},\mathbf{V},\mathbf{\Lambda},\mathbf{X};g_{n},V_{nl},\Lambda_{nl}\ and,\ X_{nl}$ correspond to the respective value at time $t$ unless specified otherwise.
[…]
$E_{\mathcal{L}}^{2}$ $=\|\nabla\mathcal{L}(\mathbf{g},\mathbf{V})-\nabla\mathcal{L}(\tilde{\mathbf{g}},\tilde{\mathbf{V}})\|^{2}$ (144)
$=\sum\limits_{n\in[N]}\left|\nabla_{g_{n}}\mathcal{L}(\mathbf{g},\mathbf{V})-\nabla_{\tilde{g}_{n}}\mathcal{L}(\tilde{\mathbf{g}},\tilde{\mathbf{V}})\right|^{2}$
$\quad+\sum\limits_{l\in[L]}\sum\limits_{n\in[N]}\left|\nabla_{V_{nl}}\mathcal{L}(\mathbf{g},\mathbf{V})-\nabla_{\tilde{V}_{nl}}\mathcal{L}(\tilde{\mathbf{g}},\tilde{\mathbf{V}})\right|^{2}.$ (145)
Substituting the gradient expressions from (25) and (31) into the above equation yields:
$E_{\mathcal{L}}^{2}$ $=\sum\limits_{n\in[N]}\left|4g_{n}\sum\limits_{l\in[L]}\Lambda_{nl}V_{nl}-4\tilde{g}_{n}\sum\limits_{l\in[L]}\tilde{\Lambda}_{nl}\tilde{V}_{nl}\right|^{2}+\sum\limits_{l\in[L]}\sum\limits_{n\in[N]}\left|2g_{n}^{2}\Lambda_{nl}-2\tilde{g}_{n}^{2}\tilde{\Lambda}_{nl}\right|^{2}.$ (146)
[…]
$E_{\mathcal{L}}^{2}$ $\leq L\sum\limits_{n\in[N]}\sum\limits_{l\in[L]}\left|4g_{n}\Lambda_{nl}V_{nl}-4\tilde{g}_{n}\tilde{\Lambda}_{nl}\tilde{V}_{nl}\right|^{2}+\sum\limits_{l\in[L]}\sum\limits_{n\in[N]}\left|2g_{n}^{2}\Lambda_{nl}-2\tilde{g}_{n}^{2}\tilde{\Lambda}_{nl}\right|^{2},$ (147)
$\leq L\sum\limits_{n\in[N]}\sum\limits_{l\in[L]}\left|4g_{n}\Lambda_{nl}V_{nl}-4\tilde{g}_{n}\tilde{\Lambda}_{nl}\tilde{V}_{nl}\right|^{2}+L\sum\limits_{l\in[L]}\sum\limits_{n\in[N]}\left|2g_{n}^{2}\Lambda_{nl}-2\tilde{g}_{n}^{2}\tilde{\Lambda}_{nl}\right|^{2}.$ (148)
Now, we try to bound the terms in each sum.
Substituting $X_{cl}\left(t\right)=g_{c}^{2}\left(t\right)V_{cl}\left(t\right)$ and using parameter bounds of domain of interest $\mathcal{D}$ in above equation yields
$\left|\tilde{X}_{cl}-X_{cl}\right|$ $=\left|\tilde{g}_{c}^{2}\tilde{V}_{cl}-g_{c}^{2}V_{cl}\right|$ (149)
$=\left|\tilde{g}_{c}^{2}\tilde{V}_{cl}-{g}_{c}^{2}\tilde{V}_{cl}+{g}_{c}^{2}\tilde{V}_{cl}-g_{c}^{2}V_{cl}\right|$ (150)
$=\left|(\tilde{g}_{c}-g_{c})(\tilde{g}_{c}+g_{c})\tilde{V}_{cl}+{g}_{c}^{2}(\tilde{V}_{cl}-V_{cl})\right|$ (151)
Substituting the parameter bounds from the domain $\mathcal{D},$
$\left|\tilde{X}_{cl}-X_{cl}\right|$ $\leq 2B_{g}B_{V}\left|\tilde{g}_{c}-g_{c}\right|+B_{g}^{2}\left|\tilde{V}_{cl}-V_{cl}\right|$ (152)
$\leq C^{2}\left(\left|\tilde{g}_{c}-g_{c}\right|+\left|\tilde{V}_{cl}-V_{cl}\right|\right),$ (153)
where, $C\geq\text{max}\{1,B_{g},2B_{V}\}$.
$|\Lambda_{nl}-\tilde{\Lambda}_{nl}|$ $=\left|\left(\mathbf{A}^{\top}\left(\mathbf{Y}-\mathbf{A}\mathbf{X}\right)\right)_{nl}-\left(\mathbf{A}^{\top}\left(\mathbf{Y}-\mathbf{A}\tilde{\mathbf{X}}\right)\right)_{nl}\right|$ (154)
$=\left|\sum\limits_{b\in[M]}\left({A}_{bn}\left({Y}_{bl}-\sum\limits_{c\in[N]}{A}_{bc}X_{cl}\right)\right)\right.-\left.\sum\limits_{b\in[M]}\left({A}_{bn}\left({Y}_{bl}-\sum\limits_{c\in[N]}{A}_{bc}\tilde{X}_{cl}\right)\right)\right|$ (155)
$=\left|\sum\limits_{b\in[M]}\left({A}_{bn}\sum\limits_{c\in[N]}{A}_{bc}\tilde{X}_{cl}\right)\right.\left.-\sum\limits_{b\in[M]}\left({A}_{bn}\sum\limits_{c\in[N]}{A}_{bc}X_{cl}\right)\right|$ (156)
$\leq\sum\limits_{c\in[N]}\left(\left|\sum\limits_{b\in[M]}{A}_{bn}{A}_{bc}\right|\left|\tilde{X}_{cl}-X_{cl}\right|\right).$ (157)
Since we have $\mathbf{A}$ is $\mu$-coherent with $\ell_{2}$-normalized columns, we have $\left|\sum\limits_{b\in[M]}{A}_{bn}{A}_{bc}\right|\leq\mu\leq 1$. Substituting, we have:
$|\Lambda_{nl}-\tilde{\Lambda}_{nl}|$ $\leq\mu\sum\limits_{c\in[N]}\left|\tilde{X}_{cl}-X_{cl}\right|.$ (158)
Substituting (153):
$|\Lambda_{nl}-\tilde{\Lambda}_{nl}|$ $\leq\mu C^{2}\sum\limits_{c\in[N]}\left(\left|\tilde{g}_{c}-g_{c}\right|+\left|\tilde{V}_{cl}-V_{cl}\right|\right).$ (159)
We can use the following looser inequality:
$|\Lambda_{nl}-\tilde{\Lambda}_{nl}|$ $\leq\mu C^{2}\sum\limits_{l\in[L]}\sum\limits_{c\in[N]}\left(\left|\tilde{g}_{c}-g_{c}\right|+\left|\tilde{V}_{cl}-V_{cl}\right|\right).$ (160)
Let $S=\sum\limits_{c\in[N]}\sum\limits_{l\in[L]}\left(\left|\tilde{g}_{c}-g_{c}\right|+\left|\tilde{V}_{cl}-V_{cl}\right|\right)$. The difference is then given by:
$|\Lambda_{nl}-\tilde{\Lambda}_{nl}|$ $\leq\mu C^{2}S.$ (161)
Similarly, we bound the term $|{g}_{n}{V}_{nl}-\tilde{g}_{n}\tilde{V}_{nl}|$ as:
$|{g}_{n}{V}_{nl}-\tilde{g}_{n}\tilde{V}_{nl}|$ $\leq|{g}_{n}-\tilde{g}_{n}\|\tilde{V}_{nl}|+|{V}_{nl}-\tilde{{V}}_{nl}\|g_{n}|.$ (162)
By applying the parameter bounds from the domain $\mathcal{D}$:
$|{g}_{n}{V}_{nl}-\tilde{g}_{n}\tilde{V}_{nl}|$ $\leq B_{V}|{g}_{n}-\tilde{g}_{n}|+B_{g}|{V}_{nl}-\tilde{{V}}_{nl}|$ (163)
$\leq\text{max}\left\{1,B_{g},B_{V}\right\}\left(\left|\tilde{g}_{n}-g_{n}\right|+\left|\tilde{V}_{nl}-V_{nl}\right|\right)$ (164)
$\leq C\left(\left|\tilde{g}_{n}-g_{n}\right|+\left|\tilde{V}_{nl}-V_{nl}\right|\right),$ (165)
where $C=\text{max}\left\{1,B_{g},2B_{V}\right\}$. Since the non-negative term, $\left|\tilde{g}_{n}-g_{n}\right|+\left|\tilde{V}_{nl}-V_{nl}\right|$, is smaller than or equal to the sum over all terms, we use the following looser upper bound for simplicity:
$|{g}_{n}{V}_{nl}-\tilde{g}_{n}\tilde{V}_{nl}|$ $\leq C\sum\limits_{l\in[L]}\sum\limits_{c\in[N]}\left(\left|\tilde{g}_{c}-g_{c}\right|+\left|\tilde{V}_{cl}-V_{cl}\right|\right)$ (166)
$|{g}_{n}{V}_{nl}-\tilde{g}_{n}\tilde{V}_{nl}|$ $\leq CS.$ (167)
Next, we find an upper bound for the magnitude of the term $|\tilde{\Lambda}_{nl}|$.
$|\tilde{\Lambda}_{nl}|$ $\leq\left|\left(\mathbf{A}^{\top}\left(\mathbf{Y}-\mathbf{A}\tilde{\mathbf{X}}\right)\right)_{nl}\right|$ (168)
$\leq\left|\sum\limits_{b\in[M]}\left({A}_{bn}\left({Y}_{bl}-\sum\limits_{c\in[N]}{A}_{bc}\tilde{X}_{cl}\right)\right)\right|.$ (169)
By applying the triangle inequality again and substituting $\tilde{X}_{cl}=\tilde{g}_{c}^{2}\tilde{V}_{cl}$:
$\leq\left|\sum\limits_{b\in[M]}{A}_{bn}{Y}_{bl}\right|+\left|\sum\limits_{b\in[M]}{A}_{bn}\sum\limits_{c\in[N]}{A}_{bc}\tilde{X}_{cl}\right|.$ (170)
Finally, applying the parameter bounds from the domain $\mathcal{D}$, and the fact that $\mathbf{A}$ is $\mu$-coherent with $\ell_{2}$-normalized columns ($|A_{ij}|\leq 1$):
$|\tilde{\Lambda}_{nl}|$ $\leq\left|\sum\limits_{b\in[M]}{A}_{bn}{Y}_{bl}\right|+\sum\limits_{c\in[N]}\left|\sum\limits_{b\in[M]}{A}_{bn}{A}_{bc}\right|\bigg|\tilde{g}_{c}^{2}\tilde{V}_{cl}\bigg|.$ (171)
Using $\bigg|\tilde{g}_{c}^{2}\tilde{V}_{cl}\bigg|\leq B_{g}^{2}B_{V}\leq C^{3}$, and $Y_{nl}\leq B_{Y}$, the above equation can be written as
$|\tilde{\Lambda}_{nl}|$ $\leq MB_{Y}+N\mu C^{3}.$ (172)
$\left|4g_{n}\Lambda_{nl}V_{nl}-4\tilde{g}_{n}\tilde{\Lambda}_{nl}\tilde{V}_{nl}\right|$ $=4|(\Lambda_{nl}-\tilde{\Lambda}_{nl}){g}_{n}{V}_{nl}+({g}_{n}{V}_{nl}-\tilde{g}_{n}\tilde{V}_{nl})\tilde{\Lambda}_{nl}|$ (173)
$\leq 4|(\Lambda_{nl}-\tilde{\Lambda}_{nl}){g}_{n}{V}_{nl}+({g}_{n}{V}_{nl}-\tilde{g}_{n}\tilde{V}_{nl})\tilde{\Lambda}_{nl}|$ (174)
$\leq 4|\Lambda_{nl}-\tilde{\Lambda}_{nl}\|{g}_{n}{V}_{nl}|+4|{g}_{n}{V}_{nl}-\tilde{g}_{n}\tilde{V}_{nl}\|\tilde{\Lambda}_{nl}|.$ (175)
Substituting $\bigg|\tilde{g}_{c}\tilde{V}_{cl}\bigg|\leq B_{g}B_{V}\leq C^{2}$, (161) and (167) in the previous equation, we have:
$\left|4g_{n}\Lambda_{nl}V_{nl}-4\tilde{g}_{n}\tilde{\Lambda}_{nl}\tilde{V}_{nl}\right|$ $\leq 4\mu C^{4}S+4(MB_{Y}+N\mu C^{3})CS$ (176)
$\leq 4\left((N+1)\mu C^{4}+MB_{Y}C\right)S.$ (177)
Next, we find an upper bound for the magnitude of the term $\left|2g_{n}^{2}\Lambda_{nl}-2\tilde{g}_{n}^{2}\tilde{\Lambda}_{nl}\right|$.
$\left|2g_{n}^{2}\Lambda_{nl}-2\tilde{g}_{n}^{2}\tilde{\Lambda}_{nl}\right|$ $\leq 2\left|g_{n}^{2}-\tilde{g}_{n}^{2}\right|\left|\tilde{\Lambda}_{nl}\right|+2\left|g_{n}^{2}\right|\left|\Lambda_{nl}-\tilde{\Lambda}_{nl}\right|$ (178)
$\leq 2\left|g_{n}-\tilde{g}_{n}\right|\left|g_{n}+\tilde{g}_{n}\right|\left|\tilde{\Lambda}_{nl}\right|+2\left|g_{n}^{2}\right|\left|\Lambda_{nl}-\tilde{\Lambda}_{nl}\right|$ (179)
$\leq 4B_{g}\left|g_{n}-\tilde{g}_{n}\right|\left|\tilde{\Lambda}_{nl}\right|+2B_{g}^{2}\left|\Lambda_{nl}-\tilde{\Lambda}_{nl}\right|$ (180)
Substituting (161) and (172) in the above equation yields:
$\left|2g_{n}^{2}\Lambda_{nl}-2\tilde{g}_{n}^{2}\tilde{\Lambda}_{nl}\right|$ $\leq 4B_{g}(MB_{Y}+N\mu C^{3})\left|g_{n}-\tilde{g}_{n}\right|+2B_{g}^{2}\mu C^{2}S.$ (181)
Note that $\left|g_{n}-\tilde{g}_{n}\right|\leq\sum\limits_{l\in[L]}\sum\limits_{c\in[N]}\left(\left|\tilde{g}_{c}-g_{c}\right|+\left|\tilde{V}_{cl}-V_{cl}\right|\right)=S$ and $B_{g}\leq C$. Substituting this in the previous equation, we have:
$\left|2g_{n}^{2}\Lambda_{nl}-2\tilde{g}_{n}^{2}\tilde{\Lambda}_{nl}\right|$ $\leq 4B_{g}(MB_{Y}+N\mu C^{3})S+2B_{g}^{2}\mu C^{2}S$ (182)
$\leq 4(C(MB_{Y}+N\mu C^{3})+\mu C^{4})S$ (183)
$\leq 4\left((N+1)\mu C^{4}+MB_{Y}C\right)S.$ (184)
Substituting the derived upper bounds ((177) and (184)) back into (148), we have:
$E_{\mathcal{L}}^{2}$ $\leq 2L\sum\limits_{n^{\prime}\in[N]}\sum\limits_{l^{\prime}\in[L]}\Big(4\left((N+1)\mu C^{4}+MB_{Y}C\right)\Big)^{2}S^{2}$ (185)
$\leq 2NL^{2}\Big(4\left((N+1)\mu C^{4}+MB_{Y}C\right)\Big)^{2}S^{2}$ (186)
$\leq 2NL^{2}\Big(4\left((N+1)\mu C^{4}+MB_{Y}C\right)\Big)^{2}\left(\sum\limits_{l\in[L]}\sum\limits_{c\in[N]}\left(\left|\tilde{g}_{c}-g_{c}\right|+\left|\tilde{V}_{cl}-V_{cl}\right|\right)\right)^{2}.$ (187)
Since $\sum\limits_{i=1}^{M}\sum\limits_{j=1}^{N}|a_{ij}|\leq\sqrt{MN\sum\limits_{i=1}^{M}\sum\limits_{j=1}^{N}|a_{ij}|^{2}}$:
$E_{\mathcal{L}}^{2}$ $\leq 4N^{2}L^{3}\Big(4\left((N+1)\mu C^{4}+MB_{Y}C\right)\Big)^{2}\sum\limits_{l\in[L]}\sum\limits_{c\in[N]}\left(\left|\tilde{g}_{c}-g_{c}\right|+\left|\tilde{V}_{cl}-V_{cl}\right|\right)^{2}$ (188)
$E_{\mathcal{L}}$ $\leq 8NL^{3/2}\left((N+1)\mu C^{4}+MB_{Y}C\right)\sqrt{\sum\limits_{l\in[L]}\sum\limits_{c\in[N]}\left(\left|\tilde{g}_{c}-g_{c}\right|+\left|\tilde{V}_{cl}-V_{cl}\right|\right)^{2}}$ (189)
$E_{\mathcal{L}}$ $\leq 16NL^{3/2}\left((N+1)\mu C^{4}+MB_{Y}C\right)\sqrt{\sum\limits_{l\in[L]}\sum\limits_{c\in[N]}\left(\left|\tilde{g}_{c}-g_{c}\right|^{2}+\left|\tilde{V}_{cl}-V_{cl}\right|^{2}\right)}.$ (190)
Thus, the loss function $\mathcal{L}(\mathbf{g},\mathbf{V})$ is $\beta$-smooth over $\mathcal{D}$ where $\beta=16NL^{3/2}\left((N+1)\mu C^{4}+MB_{Y}C\right)$ for $g_{n},V_{nl}\in\mathcal{D}\;\forall\;n\in[N],l\in[L]$. ∎
@@
Let the estimated trajectory, $\mathbf{X}(t)$ be parametrized by $({\mathbf{g}}(t),{\mathbf{V}}(t))$ and a reference trajectory $\tilde{\mathbf{X}}(t)$, be parameterized by $(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))$ both of which evolve under under continuous gradient flow, as described in Lemma C.2. Let the parameters $({\mathbf{g}}(t),{\mathbf{V}}(t))$ and $(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))$ be at a distance $D$ from the origin. At any time $t\in[0,T]$, the following inequality holds:
$\left\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\right\|^{2}_{F}$ $\leq 8D^{4}\sum\limits_{n\in[N]}\sum\limits_{l\in[L]}\left(\left({V}_{nl}(t)-\tilde{{V}}_{nl}(t)\right)^{2}\right.\left.+\left({g}_{n}(t)-\tilde{g}_{n}(t)\right)^{2}\right).$ (191)
@@
Using the inequalities $\left\{|g_{n}(t)|,\|V_{n:}(t)\|_{2}\right\}\leq\left\|\left(\mathbf{g}(t),\mathbf{V}(t)\right)\right\|\leq D$ and $\left\{\|\tilde{g}_{n}(t)|,\|\tilde{V}_{n:}(t)\|_{2}\right\}\leq\left\|\left(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t)\right)\right\|\leq D$, we can write the above equation as:
@@
$\leq 8D^{4}\sum\limits_{n\in[N]}\sum\limits_{l\in[L]}\left(\left({V}_{nl}(t)-\tilde{{V}}_{nl}(t)\right)^{2}\right.\left.+\left({g}_{n}(t)-\tilde{g}_{n}(t)\right)^{2}\right).$ (199)
C.1 Proof of Theorem V.1
@@
We begin the proof using the idealized rank-$K$ trajectory, $\tilde{\mathbf{X}}(t)$ as in (116). This initialization along with Lemmas 5.3 and B.6 ensures that the rows outside the support $\mathcal{S}_{K}$ are initialized at zero and remain zero under the gradient flow dynamics. Thus, $\tilde{\mathbf{X}}(t)$ is guaranteed to be an at-most rank-$K$ matrix for all $t\geq 0$.

### v2

@@ 2. Case II:
Let the estimated trajectory, $\mathbf{X}(t)$ be parametrized by $({\mathbf{g}}(t),{\mathbf{V}}(t))$ and a reference trajectory $\tilde{\mathbf{X}}(t)$, be parameterized by $(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))$ both of which evolve under the continuous gradient flow, as described in Lemma 5.1. Let $T>0$ and $D>0$ scalars. Specifically, assume that for all $t\in[0,T]$ and all $n\in[N]$, the condition ${\max}\left\{|g_{n}(t)|,\|V_{n:}(t)\|_{2},|\tilde{g}_{n}(t)|,\|\tilde{V}_{n:}(t),\|_{2}\right\}\leq D$ holds.
At any time $t\in[0,T]$, the following inequality holds:
$\left\|\mathbf{X}(t)-\tilde{\mathbf{X}}(t)\right\|^{2}_{F}$ $\leq 8D^{4}\Big\|(\mathbf{g}(t),\mathbf{V}(t))\left.-(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))\right\|^{2},$ (143)
where, $\Big\|(\mathbf{g}(t),\mathbf{V}(t))\left.-(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))\right\|^{2}=\sum\limits_{n\in[N]}\sum\limits_{l\in[L]}\left(\left({V}_{nl}(t)-\tilde{{V}}_{nl}(t)\right)^{2}\right.\left.+\left({g}_{n}(t)-\tilde{g}_{n}(t)\right)^{2}\right)$.
@@
Using the inequalities $\left\{|g_{n}(t)|,\|V_{n:}(t)\|_{2}\right\}\leq D$ and $\left\{\|\tilde{g}_{n}(t)|,\|\tilde{V}_{n:}(t)\|_{2}\right\}\leq D$, we can write the above equation as:
@@
$\leq 8D^{4}\sum\limits_{n\in[N]}\sum\limits_{l\in[L]}\left(\left({V}_{nl}(t)-\tilde{{V}}_{nl}(t)\right)^{2}\right.\left.+\left({g}_{n}(t)-\tilde{g}_{n}(t)\right)^{2}\right)$ (151)
Using the squared parameter distance defined in equation (82):
$\leq 8D^{4}\Big\|(\mathbf{g}(t),\mathbf{V}(t))\left.-(\tilde{\mathbf{g}}(t),\tilde{\mathbf{V}}(t))\right\|^{2}.$ (152)
C.1 Proof of Theorem 5.6
@@
A standard approach to bounding the divergence between the two trajectories is to employ Grönwall’s inequality (Razin et al., 2021). However, this technique yields an error bound that grows exponentially with time. This forces the initializations to be set far below the machine precision, rendering the theoretical guarantees practically unrealizable.
To address this, we construct a Lyapunov function, $\mathcal{E}(t)$. This approach allows us to track the divergence more tightly, leading to a bound that linearly $t$ rather than exponentially. Consequently, we derive convergence guarantees that hold for initializations that are numerically implementable.
@@
We begin the proof using the idealized rank-$K$ trajectory, $\tilde{\mathbf{X}}(t)$ as in (116). This initialization along with Corollary 5.4 and Lemma B.7 ensures that the rows outside the support $\mathcal{S}_{K}$ are initialized at zero and remain zero under the gradient flow dynamics. Thus, $\tilde{\mathbf{X}}(t)$ is guaranteed to be an at-most rank-$K$ matrix for all $t\geq 0$.